OpenStudy (anonymous):

Find all the zeros of the equation. -4x^4 - 44x^2 + 3600 = 0 A. 5, -5, 6i, -6i B. 5, 6i C. 5, -5, 6i, 0 D. -5, -6i

OpenStudy (anonymous):

\[-4x^4 - 44x^2 + 3600 = 0\] divide all by \(-4\) get \[x^2+11x-900=0\]

OpenStudy (anonymous):

sorry i mean get \[x^4+11x^2-900=0\]

OpenStudy (anonymous):

by some miracle this one actually factors as \[(x^2-25)(x^2+36)=0\] and it should be good to go from there

OpenStudy (anonymous):

actually now that i look at it this problem required no work at all 'it is a 4th degree polynomial so it must have 4 zeros while it is true that some could repeat, if \(6i\) is a zero, then its conjugate \(-6i\) must also be one so without any computation we know A is the only possible answer

OpenStudy (anonymous):

good to remember if you have to take a test or quiz if \(a+bi\) is a zero of a polynomial with real coefficients, then so is \(a-bi\) its conjugate